# ManiCM: Real-time 3D Diffusion Policy via Consistency Model for Robotic Manipulation

### v2

@@ I Introduction
However, diffusion-based fashions inevitably suffer from severe runtime inefficiency by requiring considerable sampling steps for high-quality action synthesis during inference, even with some sampling acceleration [24]. As shown in Figure 1, the diffusion policy (DP [5]) requires 162ms per step to generate a high-quality action. Moreover, the 3D diffusion policy (DP3 [6]) only achieves a decision latency of 178ms when dealing with the high-dimensional 3D point cloud input, which blocks the applications of real-time closed-loop control. To address this issue, recent efforts have been proposed to speed up the inference phase of diffusion models by hierarchical sampling [25, 15, 16]. Nevertheless, it is hard to determine the hierarchy across task domains that exhibit different challenging levels, which limits the practicality of the existing works to 3D robotic manipulation.
Recently, the concept of consistency models [26, 27] has been introduced in image generation, which synthetizes high-fidelity images with a minimal number of steps. In this paper, we propose a ManiCM method that leverages the concept of consistency models for real-time robotic manipulation. Different from existing methods that focus on designing hierarchical planning for specific domains, we design a generalizable consistency distillation technique to generate robot actions in only one-step inference. More specifically, we first formulate a diffusion process to denoise robot action conditioned on the point cloud, and then ensure the consistency property that the original action can be directly denoised from any point along the ODE trajectory. Then, we utilize the consistency distillation technique to predict the action sample directly, based on the observation that predicting the denoised action converges faster than predicting the noise in the vision community. We evaluate ManiCM on a diverse set of robotic manipulation tasks from Adroit and Metaworld, and the results illustrate that our approach achieves an average inference speed acceleration of 10 compared to the state-of-the-art method, while still maintaining competitive average success rates (Figure 1). We introduce consistency distillation into the robotic manipulation area and accelerate high-quality 3D manipulation action generation to a real-time level. Our contributions can be concluded as follows:
@@ III Approach
In this section, we first present the problem formulation and the preliminaries on consistency models (Section III-A). Then, we present an overview of our pipeline, and detail the design of each component in the manipulation consistency model (Section III-B). To enable our model to learn consistent generation from the teacher network, we propose the manipulation consistency distillation (Section III-C).
@@ III-A Problem Formulation
As shown in Equation 1, the implication of the self-consistency property is that for any input pairs on the same PF-ODE trajectory, their outputs remain consistent. All consistency models have to meet the boundary condition , as it plays an important role in the success of consistency model training, according to [26]. To impose the self-consistency property on deep neural networks, we implement function as:
@@
where is a metric function chosen for measuring the distance between two samples. and are referred to as ‘online network’ and ‘target network’ according to [26]. When the optimization of the online network converges, the target network will eventually match the online network since is a running average of .
@@ III-B ManiCM: Manipulation Consistency Model
Manipulation Self-Consistency Function. However, the iterative procedure in Equation 4 raises a non-negligible computational burden, especially with high-dimensional point cloud conditions. To this end, we propose to constrain the original action to be recovered from any point along the trajectory, so that the model can generate action in only one step in the test phase:
where is a manipulation consistency function that contains the accumulated in Equation 4 along the trajectory. In practice, , , where denotes a timestep scaling value while is a balance value. The can be parametrized to predict the action sample or the noise epsilon. In mainstream diffusion generation works for image generation, the default prediction type is noise rather than sample. We speculate that while directly predicting samples may intuitively yield higher-quality generated outputs [43], predicting the distribution of images is much more challenging for denoising models in high-dimensional image space. However, the dimensionality of the robot action is quite small (approximately ) in manipulation tasks. In this scenario, directly learning the distribution of the samples is relatively easy and brings low variance and fast convergence, echoing [6]. More importantly, the large prediction variance caused by predicting noise will be amplified in consistency models. Specifically, the consistency property encourages the to match , while the original diffusion model only requires to be close to . Thus, the variance is propagated along the ODE trajectory in consistency modeling, which ultimately blocks the convergence of prediction noise in manipulation tasks (Figure 3). Therefore, we parameterize the manipulation self-consistency function to predict the action sample rather than noise.
@@ III-C Manipulation Consistency Distillation
To train ManiCM, we adopt the method of consistency distillation, as depicted in Figure 2. Here represents the sequence of robotic arm actions, which comprises multiple actions, with the dimensionality determined by the specific task. denotes the point cloud obtained from images while signifies the current pose of the robotic arm from the offline dataset. These inputs are processed through an encoder to obtain compact representations, which are then used as conditional inputs for the teacher network, online network, and target network. Given a ground-truth expert action , we first obtain the noisy action by conducting a forward diffusion operation with steps to add noise to , where is the skipping interval mentioned in Section III-A. Subsequently, the noisy action is feed-forwarded to a frozen teacher network and the online network to decode the clean and , respectively. In order to make sure the self-consistency property Equation 1, we design a target network cloned from the online network, and we expect the denoised outputs of the online network and the target network to be aligned. Specifically, the target network generates utilizing obtained from processing with a -step ODE solver (e.g., DDIM [24]), such that and are obtained by the consistency function in Equation 2. Therefore, the consistency distillation loss is expressed as,
where is the norm. Besides, is updated with the exponential moving average (EMA) of the parameter defined as , where denotes the stopgrad operation and satisfies . In Equation 6, is the -step estimation from . Therefore, the can be formulated as,
where is the update function of a -step ODE solver applied to PF-ODE. As outlined in Section III-A, we utilize EMA to update the parameters of the target network from the parameters of the online network. We refer to the pre-trained manipulation 3D diffusion model, such as DP3 [6], as the "teacher network".
Compared to other manipulation diffusion models, our ManiCM model achieves the fastest runtime (16ms per action) during the inference phase by sampling high-quality actions in just one step, which is illustrated in Table I.
@@ IV-A Experimental Setup
Evaluation Metrics. Following [6], we assess episodes every training epochs and then calculate the average of the highest success rates, along with the average runtime per step. For each main experiment, we run random seeds with seed numbers , , to mitigate the performance fluctuation. We also report the Number of Function Evaluations (NFE) as an additional device-agnostic metric for inference speed. For the success rate and runtime, the higher the better. For the NFE, the lower the better.
Implementation Details. For ManiCM, an AdamW [49] optimizer with a batch size of 128 and a learning rate of 5e-5 is employed for 3K epochs training. We also adopt a cosine decay learning rate scheduler and 500 iterations of linear warm-up. The EMA rate is set to by default and we use DDIM-Solver [24] with a skipping step of as ODE Solvers. Unless otherwise specified, our experiments are conducted using training time steps. The teacher model employed in our experiments uses DDIM with an inference step of . For the input, We utilize an observation steps value of , indicating that we use the point clouds from the two most recent time steps as conditions. Similarly, with a horizon of , we predict four consecutive actions during each iteration for long-horizon tasks, while only actions are executed per prediction. To focus on the interest area, the image size is cropped to pixels. The point cloud is downsampled to points using the Farthest Point Sampling (FPS) algorithm, while the dimension of agent poses and actions are determined by the specific task. Both point clouds and agent poses are encoded using an MLP encoder into embeddings of length . For normalization, we limit the min and max of action and observation to independently since the prediction of DDIM would clip the prediction to to ensure training stability. We implement our model in PyTorch [50], and all compared models are trained and evaluated on one NVIDIA RTX 4090 GPU to benchmark the decision runtime.
@@ IV-B Comparison with the State-of-the-Art Methods
Comparisons on Runtime. Table I illustrates the comparison of the runtime per step of each model. Our method achieves an impressive runtime around 16ms per step on average, which is 10 faster than the baseline method DP3. The acceleration comes from the manipulation consistency model, which can generate robot actions in only one-step inference. Besides, the -step inference version of ManiCM also improves the efficiency by 3, Moreover, the proposed ManiCM also surpasses the lightweight version Simple DP3 significantly, which shows that the speedup from reducing the sampling speed is a more efficient way than shrinking the model itself. The results demonstrate the superiority of our ManiCM in terms of runtime, which verifies our intuition that incorporating the consistency model boosts the inference speed in various robotic manipulation tasks.
Comparisons on Success Rate. Table II depicts the comparison of the success of each model. The results show that our method accelerates the diffusion process without any performance drop. Specifically, with only one-step inference, our ManiCM can approximate or even surpass the state-of-the-art model DP3 and Simple DP3 by %. By distilling the diffusion process with consistency regularization, our method avoids overfitting to the training demonstrations. Especially, in the Metaworld Very Hard tasks, our method outperforms the state-of-the-art method with an absolute improvement of %. Moreover, increasing the sampling steps in ManiCM yields even better performance on average (% vs. %). The above experimental results prove the effectiveness of handling robotic tasks with our manipulation consistency model.
Comparisons on Multi-Task Setup. Figure 5 compares ManiCM with the 3D Diffuser Actor [29] baseline on the multi-task RLBbench simulation environment. The results show that ManiCM achieves competitive performance while demonstrating a remarkable 13.3× acceleration in inference speed, highlighting our consistency model’s effectiveness on a multi-task setup.
@@ IV-C Real World Experiments
As shown in Figure 6, we evaluate ManiCM in the real world on three tasks from easy to hard with different embodiments: Slide Block, Stack Cube, and Transfer Piece.
@@
We use an end effector and a DH Robotics AG-95 gripper based on UR3e arms and a Realsense D435i camera to capture visual observations, with an NVIDIA RTX 4090 GPU for inference. We retain the same network architecture and input-output configuration as in the simulation environment. For each task, we conduct 10 trials to evaluate the average success rate and inference time. As demonstrated in Table IV, ManiCM achieves 10× faster inference speeds than DP3 while maintaining comparable success rates, validating its effectiveness in real-world setup.
@@ IV-D Ablation Studies
Ablation on Sample Prediction and Epsilon Prediction. We conduct a comparison between directly predicting the sample and predicting the noise ‘epsilon’, which is a widely-used technique in the vision community [28]. From Figure 3, we can observe that sample prediction in the noise sampler brings faster convergence than predicting the noise epsilon in various tasks and methods. This is also observed in [6]. Besides, the performance of the noise prediction is even worse when we implement consistency distillation. This is because predicting the action sample is more direct than training a denoising model for the low-dimensional action manifold. More importantly, the consistency distillation process will cause compound instability, so that the performance of sample prediction is much stronger than noise prediction variation (% vs. % on average).
Ablation on Different Diffusion Steps of the Teacher Model. We compare our ManiCM with the state-of-the-art method DP3 under different DDIM sampling steps in Table III. We can conclude that: 1) The task success rates do not monotonically increase when taking more diffusion sampling steps in both DP3 and ManiCM, which indicates that most of the diffusion steps are redundant. As a result, we can distill the teacher model with multiple-step inference into an online model that generates high-quality action in only one sampling step. 2) For different diffusion sampling steps, our ManiCM can stably distill competitive capacity in task completion, where it accelerates the runtime efficiency by 57 when the teacher is a DDIM with inference steps.
@@ IV-E Qualitative Analysis
We present two qualitative examples of the generated action sequence in Figure 4 from DP3 and our ManiCM in Adroit and MetaWorld, respectively. In the left case, the agent needs to spin the pen to a specific angle given by the green reference pen with its dexterous hand. The compared method DP3 fails to spin the pen to the correct angle, while our method succeeds in spinning it to the angle specified by the green pen on the ground. In the right case, the agent is required to kick the football into the goal with its parallel jaw gripper. DP3 stretches the gripper to the goal without holding the ball by just mimicking the expert trajectories in the training set, while ManiCM successfully kicks the ball into the goal. Both cases verify that our ManiCM accelerates the denoising process by a large margin, while maintaining the fine-grained 3D comprehension across diverse robotic tasks.

### v3

@@ I Introduction
However, diffusion-based fashions inevitably suffer from severe runtime inefficiency by requiring considerable sampling steps for high-quality action synthesis during inference, even with some sampling acceleration [24]. As shown in Figure˜1, the diffusion policy (DP [5]) requires 162ms per step to generate a high-quality action. Moreover, the 3D diffusion policy (DP3 [6]) only achieves a decision latency of 178ms when dealing with the high-dimensional 3D point cloud input, which blocks the applications of real-time closed-loop control. To address this issue, recent efforts have been proposed to speed up the inference phase of diffusion models by hierarchical sampling [25, 15, 16]. Nevertheless, it is hard to determine the hierarchy across task domains that exhibit different challenging levels, which limits the practicality of the existing works to 3D robotic manipulation.
Recently, the concept of consistency models [26, 27] has been introduced in image generation, which synthetizes high-fidelity images with a minimal number of steps. In this paper, we propose a ManiCM method that leverages the concept of consistency models for real-time robotic manipulation. Different from existing methods that focus on designing hierarchical planning for specific domains, we design a generalizable consistency distillation technique to generate robot actions in only one-step inference. More specifically, we first formulate a diffusion process to denoise robot action conditioned on the point cloud, and then ensure the consistency property that the original action can be directly denoised from any point along the ODE trajectory. Then, we utilize the consistency distillation technique to predict the action sample directly, based on the observation that predicting the denoised action converges faster than predicting the noise in the vision community. We evaluate ManiCM on a diverse set of robotic manipulation tasks from Adroit and Metaworld, and the results illustrate that our approach achieves an average inference speed acceleration of 10 compared to the state-of-the-art method, while still maintaining competitive average success rates (Figure˜1). We introduce consistency distillation into the robotic manipulation area and accelerate high-quality 3D manipulation action generation to a real-time level. Our contributions can be concluded as follows:
@@ III Approach
In this section, we first present the problem formulation and the preliminaries on consistency models (Section˜III-A). Then, we present an overview of our pipeline, and detail the design of each component in the manipulation consistency model (Section˜III-B). To enable our model to learn consistent generation from the teacher network, we propose the manipulation consistency distillation (Section˜III-C).
@@ III-A Problem Formulation
As shown in Equation˜1, the implication of the self-consistency property is that for any input pairs on the same PF-ODE trajectory, their outputs remain consistent. All consistency models have to meet the boundary condition , as it plays an important role in the success of consistency model training, according to [26]. To impose the self-consistency property on deep neural networks, we implement function as:
@@
where is a metric function chosen for measuring the distance between two samples and represent the ODE Solver. and are referred to as ‘online network’ and ‘target network’ according to [26]. When the optimization of the online network converges, the target network will eventually match the online network since is a running average of .
@@ III-B ManiCM: Manipulation Consistency Model
Manipulation Self-Consistency Function. However, the iterative procedure in Equation˜4 raises a non-negligible computational burden, especially with high-dimensional point cloud conditions. To this end, we propose to constrain the original action to be recovered from any point along the trajectory, so that the model can generate action in only one step in the test phase:
where is a manipulation consistency function that contains the accumulated in Equation˜4 along the trajectory. In practice, , , where denotes a timestep scaling value while is a balance value. The can be parametrized to predict the action sample or the noise epsilon. In mainstream diffusion generation works for image generation, the default prediction type is noise rather than sample. We speculate that while directly predicting samples may intuitively yield higher-quality generated outputs [43], predicting the distribution of images is much more challenging for denoising models in high-dimensional image space. However, the dimensionality of the robot action is quite small (approximately ) in manipulation tasks. In this scenario, directly learning the distribution of the samples is relatively easy and brings low variance and fast convergence, echoing [6]. More importantly, the large prediction variance caused by predicting noise will be amplified in consistency models. Specifically, the consistency property encourages the to match , while the original diffusion model only requires to be close to . Thus, the variance is propagated along the ODE trajectory in consistency modeling, which ultimately blocks the convergence of prediction noise in manipulation tasks (Figure˜3). Therefore, we parameterize the manipulation self-consistency function to predict the action sample rather than noise.
@@ III-C Manipulation Consistency Distillation
To train ManiCM, we adopt the method of consistency distillation, as depicted in Figure˜2. Here represents the sequence of robotic arm actions, which comprises multiple actions, with the dimensionality determined by the specific task. denotes the point cloud obtained from images while signifies the current pose of the robotic arm from the offline dataset. These inputs are processed through an encoder to obtain compact representations, which are then used as conditional inputs for the teacher network, online network, and target network. Given a ground-truth expert action , we first obtain the noisy action by conducting a forward diffusion operation with steps to add noise to , where is the skipping interval mentioned in Section˜III-A. Subsequently, the noisy action is feed-forwarded to a frozen teacher network and the online network to decode the clean and , respectively. In order to make sure the self-consistency property Equation˜1 is respected, we design a target network cloned from the online network, and we expect the denoised outputs of the online network and the target network to be aligned. Specifically, the target network generates utilizing obtained from processing with a -step ODE solver (e.g., DDIM [24]), such that and are obtained by the consistency function in Equation˜2. Therefore, the consistency distillation loss is expressed as,
where is the norm. Besides, is updated with the exponential moving average (EMA) of the parameter defined as , where denotes the stopgrad operation and satisfies . In Equation˜6, is the -step estimation from . Therefore, the can be formulated as,
where is the update function of a -step ODE solver applied to PF-ODE. As outlined in Section˜III-A, we utilize EMA to update the parameters of the target network from the parameters of the online network. We refer to the pre-trained manipulation 3D diffusion model, such as DP3 [6], as the "teacher network".
Compared to other manipulation diffusion models, our ManiCM model achieves the fastest runtime (16ms per action) during the inference phase by sampling high-quality actions in just one step, which is illustrated in Table˜I.
@@ IV-A Experimental Setup
Evaluation Metrics. Following [6], we assess episodes every training epochs and then calculate the average of the highest success rates, along with the average runtime per step. For each main experiment, we run random seeds with seed numbers , , to mitigate the performance fluctuation. We also report the Number of Function Evaluations (NFE) as an additional device-agnostic metric for inference speed. For the success rate, the higher the better. For the NFE and the runtime, the lower the better.
Implementation Details. For ManiCM, an AdamW [49] optimizer with a batch size of 128 and a learning rate of 5e-5 is employed for 3K epochs training. We also adopt a cosine decay learning rate scheduler and 500 iterations of linear warm-up. The EMA rate is set to by default and we use DDIM-Solver [24] with a skipping step of as ODE Solvers. Unless otherwise specified, our experiments are conducted using training time steps. The teacher model employed in our experiments uses DDIM with an inference step of . For the input, We utilize an observation steps value of , indicating that we use the point clouds from the two most recent time steps as conditions. Similarly, with a horizon of , we predict four consecutive actions during each iteration for long-horizon tasks, while only actions are executed per prediction. To focus on the interest area, we obtain depth images with size 84 × 84 from a single camera. The point cloud is downsampled to points using the Farthest Point Sampling (FPS) algorithm, while the dimension of agent poses and actions are determined by the specific task. Both point clouds and agent poses are encoded using an MLP encoder into embeddings of length . For normalization, we limit the min and max of action and observation to independently since the prediction of DDIM would clip the prediction to to ensure training stability. We implement our model in PyTorch [50], and all compared models are trained and evaluated on one NVIDIA RTX 4090 GPU to benchmark the decision runtime.
@@ IV-B Comparison with the State-of-the-Art Methods
Comparisons on Runtime. Table˜I illustrates the comparison of the runtime per step of each model. Our method achieves an impressive runtime around 16ms per step on average, which is 10 faster than the baseline method DP3. The acceleration comes from the manipulation consistency model, which can generate robot actions in only one-step inference. Besides, the -step inference version of ManiCM also improves the efficiency by 3, Moreover, the proposed ManiCM also surpasses the lightweight version Simple DP3 significantly, which shows that the speedup from reducing the sampling speed is a more efficient way than shrinking the model itself. The results demonstrate the superiority of our ManiCM in terms of runtime, which verifies our intuition that incorporating the consistency model boosts the inference speed in various robotic manipulation tasks.
Comparisons on Success Rate. Table˜II depicts the comparison of the success of each model. The results show that our method accelerates the diffusion process without any performance drop. Specifically, with only one-step inference, our ManiCM can approximate or even surpass the state-of-the-art model DP3 and Simple DP3 by %. By distilling the diffusion process with consistency regularization, our method avoids overfitting to the training demonstrations. Especially, in the Metaworld Very Hard tasks, our method outperforms the state-of-the-art method with an absolute improvement of %. Moreover, increasing the sampling steps in ManiCM yields even better performance on average (% vs. %). The above experimental results prove the effectiveness of handling robotic tasks with our manipulation consistency model.
Comparisons on Multi-Task Setup. Figure˜5 compares ManiCM with the 3D Diffuser Actor [29] baseline on the multi-task RLBbench simulation environment. The results show that ManiCM achieves competitive performance while demonstrating a remarkable 13.3× acceleration in inference speed, highlighting our consistency model’s effectiveness on a multi-task setup.
@@ IV-C Real World Experiments
As shown in Figure˜6, we evaluate ManiCM in the real world on three tasks from easy to hard with different embodiments: Slide Block, Stack Cube, and Transfer Piece.
@@
We use an end effector and a DH Robotics AG-95 gripper based on UR3e arms and a Realsense D435i camera to capture visual observations, with an NVIDIA RTX 4090 GPU for inference. We retain the same network architecture and input-output configuration as in the simulation environment. For each task, we conduct 10 trials to evaluate the average success rate and inference time. As demonstrated in Table˜IV, ManiCM achieves 10× faster inference speeds than DP3 while maintaining comparable success rates, validating its effectiveness in real-world setup.
@@ IV-D Ablation Studies
Ablation on Sample Prediction and Epsilon Prediction. We conduct a comparison between directly predicting the sample and predicting the noise ‘epsilon’, which is a widely-used technique in the vision community [28]. From Figure˜3, we can observe that sample prediction in the noise sampler brings faster convergence than predicting the noise epsilon in various tasks and methods. This is also observed in [6]. Besides, the performance of the noise prediction is even worse when we implement consistency distillation. This is because predicting the action sample is more direct than training a denoising model for the low-dimensional action manifold. More importantly, the consistency distillation process will cause compound instability, so that the performance of sample prediction is much stronger than noise prediction variation (% vs. % on average).
Ablation on Different Diffusion Steps of the Teacher Model. We compare our ManiCM with the state-of-the-art method DP3 under different DDIM sampling steps in Table˜III. We can conclude that: 1) The task success rates do not monotonically increase when taking more diffusion sampling steps in both DP3 and ManiCM, which indicates that most of the diffusion steps are redundant. As a result, we can distill the teacher model with multiple-step inference into an online model that generates high-quality action in only one sampling step. 2) For different diffusion sampling steps, our ManiCM can stably distill competitive capacity in task completion, where it accelerates the runtime efficiency by 57 when the teacher is a DDIM with inference steps.
@@ IV-E Qualitative Analysis
We present two qualitative examples of the generated action sequence in Figure˜4 from DP3 and our ManiCM in Adroit and MetaWorld, respectively. In the left case, the agent needs to spin the pen to a specific angle given by the green reference pen with its dexterous hand. The compared method DP3 fails to spin the pen to the correct angle, while our method succeeds in spinning it to the angle specified by the green pen on the ground. In the right case, the agent is required to kick the football into the goal with its parallel jaw gripper. DP3 stretches the gripper to the goal without holding the ball by just mimicking the expert trajectories in the training set, while ManiCM successfully kicks the ball into the goal. Both cases verify that our ManiCM accelerates the denoising process by a large margin, while maintaining the fine-grained 3D comprehension across diverse robotic tasks.